hey. i'm pretty good at radicals, but i'm completely stumped on this one.
6√24x^5
______
18x
i'd appreciate your help.

Answer by fractalier(6550)

Okay...starting from
6(rad(24x^5)) / 18x
Let us pull the perfect squares out of the 24x^5 and we get
6(rad(24x^5)) / 18x =
6(rad(4x^4))(rad(6x)) / 18x =
6(2x^2)(rad(6x)) / 18x =
(12x^2)(rad(6x)) / 18x
Now do some cancelling...
2x(rad(6x)) / 3